1.     A stone is thrown vertically upward from ground level. Six seconds later, a second stone is also thrown vertically upward from the same point with a velocity of 40 m/sec. If the two stones meet 30 meters above the ground with the first stone already going down, with what velocity was the first stone thrown?



1
Expert's answer
2022-03-26T15:30:49-0400

"v_2=gt-v_1+\sqrt{2gh}=20~\frac ms."